# A Hessian for Gaussian Mixture Likelihoods in Nonlinear Least Squares

### v1

@@ VII-C Lost in the Woods Dataset
${}^{\dagger}$ 18.7 22.3 27.0 16.7 18.4 24.2
NLS ${}^{\dagger}$ 18.7 22.2 26.9 16.6 18.4 23.9
The “Lost in the Woods” dataset [15] consists of a wheeled robot navigating a forest of plastic tubes, which are treated as landmarks. The robot receives wheel odometry measurements providing forward velocity and angular velocity measurements. The robot is equipped with a laser rangefinder that provides range-bearing measurements to the landmarks. The task is to estimate the robot poses, $\mathcal{T}=\left\{\mathbf{T}_{1}^{\mathrm{rob}},\ldots,\mathbf{T}_{K}^{\mathrm{rob}}\right\}$, with $\mathbf{T}_{i}^{\mathrm{rob}}\in SE(2),i=1,\ldots,K$, and the landmark positions =$\mathcal{L}=\left\{\boldsymbol{\ell}_{1},\ldots,\boldsymbol{\ell}_{n_{\ell}}\right\},\boldsymbol{\ell}_{i}\in\mathbb{R}^{2},i=1,\ldots,n_{\ell}$, given the odometry and range-bearing measurements. The process model consists of nonholonomic vehicle kinematics with the wheel odometry forward and angular velocity inputs detailed in [22]. The range-bearing measurements of landmark $j$ observed from pose $k$ are in the form $\mbf{y}_{jk}=\mbf{g}\left(\mbf{T}_{k}^{\mathrm{rob}},\boldsymbol{\ell}_{j}\right)+\mbf{v}_{jk}$, where $\mbf{v}_{jk}\sim\mathcal{N}\left(\mbf{0},\mbf{R}_{jk}\right)$ and $\mbf{g}\left(\cdot,\cdot\right)$ is the range-bearing measurement model, also detailed in [22]. While the dataset provides the data-association variables (i.e., which landmark each measurement corresponds to), to evaluate the performance of the algorithms, the challenging example of unknown data associations is considered. When the data associations are unknown, a multimodal measurement likelihood may be used, as in [4], where each measurement could have been produced from any landmark in $\mathcal{L}$, such that
@@
The optimization problem to be solved then consists of process model residuals corresponding to the nonholonomic vehicle dynamics, as well as multimodal range-bearing error terms of the form (48) for each measurement. To simplify the problem, it is assumed that the groundtruth number of landmarks are known, so that the number of components in each mixture (48) is known a priori. Additionally, it is assumed that an initial guess for each landmark position is available. This is similar to the setup of the “oracle baseline” described in [23]. The dataset provides odometry and landmark measurements at 10 Hz. To reduce the problem size, the landmark measurements are subsampled to 1 Hz.
@@
For each subsequence length and maximum range radius, the average RMSE and the average number of iterations are summarized in Table III and Table IV, respectively. The difficulty of the problem increases with the subsequence length, because the deadreckoned initialization becomes further from the true solution. The difficulty also increases with a smaller maximum measurement radius, since fewer measurements are used. For all methods, the RMSE increases with the scenario difficulty, as expected. The HSM method provides a performance improvement in the $r_{\textrm{max}}=2\ \text{m},T_{\text{subseq}}=40\ \text{s}$ case, which corresponds to the hardest difficulty considered. In the other scenarios, all methods perform similarly. The number of iterations also presents a slight improvement in the same case of $r_{\textrm{max}}=2\ \text{m},T_{\text{subseq}}=40\ \text{s}$, while showing similar performance in the other scenarios.

### v2

@@ VII-C Lost in the Woods Dataset
HSM 18.7 22.3 27.0 16.7 18.4 23.9
TABLE V: Average time on the Lost in the Woods subsequences.
Max Radius (m) 2 4
Subseq. Len (s) 20 30 40 20 30 40
MM 1.45 3.08 5.22 2.88 5.55 10.46
SM 15.04 35.27 49.27 48.86 87.48 121.64
MSM 1.63 3.53 6.22 3.57 7.16 13.76
HSM 1.60 3.41 5.83 3.38 6.66 12.76
The “Lost in the Woods” dataset [lostInTheWoodsDataset] consists of a wheeled robot navigating a forest of plastic tubes, which are treated as landmarks. The robot receives wheel odometry measurements providing forward velocity and angular velocity measurements. The robot is equipped with a laser rangefinder that provides range-bearing measurements to the landmarks. The task is to estimate the robot poses, $\mathcal{T}=\left\{\mathbf{T}_{1}^{\mathrm{rob}},\ldots,\mathbf{T}_{K}^{\mathrm{rob}}\right\}$, with $\mathbf{T}_{i}^{\mathrm{rob}}\in SE(2),i=1,\ldots,K$, and the landmark positions =$\mathcal{L}=\left\{\boldsymbol{\ell}_{1},\ldots,\boldsymbol{\ell}_{n_{\ell}}\right\},\boldsymbol{\ell}_{i}\in\mathbb{R}^{2},i=1,\ldots,n_{\ell}$, given the odometry and range-bearing measurements. The process model consists of nonholonomic vehicle kinematics with the wheel odometry forward and angular velocity inputs detailed in [cossette2020]. The range-bearing measurements of landmark $j$ observed from pose $k$ are in the form $\mbf{y}_{jk}=\mbf{g}\left(\mbf{T}_{k}^{\mathrm{rob}},\boldsymbol{\ell}_{j}\right)+\mbf{v}_{jk}$, where $\mbf{v}_{jk}\sim\mathcal{N}\left(\mbf{0},\mbf{R}_{jk}\right)$ and $\mbf{g}\left(\cdot,\cdot\right)$ is the range-bearing measurement model, also detailed in [cossette2020]. While the dataset provides the data-association variables (i.e., which landmark each measurement corresponds to), to evaluate the performance of the algorithms, the challenging example of unknown data associations is considered. When the data associations are unknown, a multimodal measurement likelihood may be used, as in [doherty2020probabilistic], where each measurement could have been produced from any landmark in $\mathcal{L}$, such that
@@
The optimization problem to be solved then consists of process model residuals corresponding to the nonholonomic vehicle dynamics, as well as multimodal range-bearing error terms of the form (43) for each measurement. To simplify the problem, it is assumed that the groundtruth number of landmarks are known, so that the number of components in each mixture (43) is known a priori. Additionally, it is assumed that an initial guess for each landmark position is available. This is similar to the setup of the “oracle baseline” described in [zhang2023]. The dataset provides odometry and landmark measurements at 10 Hz. To reduce the problem size, the landmark measurements are subsampled to 1 Hz.
@@
For each subsequence length and maximum range radius, the average RMSE and the average number of iterations are summarized in Table III and Table IV, respectively. With the exception of Sum-Mixture that fails due to the inacurate Hessian, performance is similar in this experiment for all methods in terms of RMSE and iteration count. This is explained by landmarks being far to each other relative to the range-bearing sensor noise characteristics. The less the components overlap, the smaller performance improvement can be expected from HSM relative to previous methods as stated in Sec. I. There is a slight improvement for runtime for HSM compared to MSM in Table V. This difference must be interpreted with caution. Although reasonable efforts were made to optimize all of the approaches, the Python implementation may still not be representative of optimized real-time systems. Nevertheless, both methods require the computation of $\frac{\partial\rho}{\partial f_{k}},\mbf{e}_{k},\frac{\partial\mbf{e}_{k}}{\partial\mbf{x}}$ for each component to compute error Jacobians at each iteration. Thus, it is recommended to consider the iteration count as an implementation-independent proxy for runtime.